OpenStudy (anonymous):

how can u find the limit for (sin2x)/x as x approaches 0?

sam (.sam.):

L'Hôpital's rule

OpenStudy (anonymous):

don't know what that is im in trig

OpenStudy (anonymous):

i learned it by finding the sin of 2x in degrees first

sam (.sam.):

\[\lim_{x \rightarrow 0} \frac{\sin(2x)}{x}\] Differentiate numerator and denominator \[ \lim_{x \rightarrow 0} \frac{2\cos(2x)}{1}\] As x approaches zero, \[= \lim_{x \rightarrow 0} 2\cos(0) \\ \\ =\lim_{x \rightarrow 0} 2(1) \\ \\ =2\]

OpenStudy (anonymous):

where did you get cosine from?

sam (.sam.):

You differentiate sin(2x) you'll get 2cos(2x)

OpenStudy (anonymous):

Without using L Hopital's Rule, you want to use the fact that:\[\lim_{u\rightarrow 0}\frac{\sin(u)}{u}=1\] Does this look familiar?

OpenStudy (anonymous):

nope i was taught to find the degree value of sin 2x first? idk .....